### v1

@@ 3.2 Deep Learning Models for Time Series Forecasting
Time series forecasting, which leverages historical data to predict future trends, has emerged as a crucial area of research with significant applications in public health. In recent years, there has been a rapid expansion in data-driven approaches for epidemic forecasting, aimed at predicting the spread and severity of infectious diseases [34, 35, 36]. These forecasts are essential for developing efficient public health surveillance systems. They enable early detection of abnormal infection patterns and adverse health outcomes while mitigating the socioeconomic impacts of epidemics. Recently, deep learning-based approaches have emerged as a more flexible tool to forecast epidemics. These methods can process complex datasets, learn intricate patterns, and capture non-linear relationships more effectively than classical statistical models for time series forecasting [15]. Encoder-decoder frameworks incorporating attention mechanisms [37] or convolutional operations [38] represent prominent approaches. Transformers, a popular example of encoder-decoder architectures with attention mechanisms, process time series data by mapping it to a high-dimensional vector through the encoder’s input layer [39]. Positional encoding adds sequential information to the data, which is then passed through multiple encoder layers equipped with self-attention and feed-forward components. The output of the encoder is fed into the decoder, which processes it through corresponding input and decoder layers. Each decoder layer applies self-attention to the input and integrates this information to produce predictions. The final output layer employs look-ahead masking and one-position offsets to ensure that forecasts rely solely on historical data points. Another encoder-decoder-based forecasting method is the Temporal Convolutional Network (TCN), which uses convolutional networks to model sequential data [40]. The encoder block in TCN architectures consists of stacked dilated causal convolutional networks, enhanced with residual blocks to capture stochastic patterns in historical time series observations. The decoder block incorporates latent representations from the encoder and generates an output through its residual network module. A dense layer further processes this output to predict the future trajectory of the time series. On the other hand, attention-free and convolution-free models with deep stacks of fully connected layers are also designed to capture input-output relationships while implicitly prioritizing resource-efficient computations. Among these models, the Neural Basis Expansion for Time Series (NBeats) framework employs a block-based architecture to forecast time series data through residual learning [41]. This framework consists of stacked blocks, with the initial block modeling the input time series to estimate the forward and backward outputs. Subsequent blocks refine the residual outputs of their predecessors similarly. Each block features a dense layer with ReLU activation, which generates forward and backward predictors of expansion coefficients. These coefficients are then processed by backward and forward basis layers to produce the desired backcast and forecast. Building on the block-based design of NBeats, the Neural Hierarchical Interpolation for Time Series (NHits) framework incorporates a multi-rate signal sampling approach to capture the multi-scale characteristics of time series data in long-term forecasting tasks [42]. This architecture includes stacked blocks, each containing a Multilayer Perceptron (MLP) and a MaxPool layer, which focus on analyzing low-frequency components.
@@
The architectural designs of these scalable deep learning forecasters demonstrate superior performance in capturing long-term dependencies compared to standard RNN-based frameworks, making them highly effective for epidemic forecasting tasks. For instance, Transformers have been shown to accurately forecast influenza cases [39], while attention mechanisms have been utilized for predicting dengue incidence in Vietnam [43]. Similarly, a modified NBeats framework was developed to forecast COVID-19 hospitalizations in Canada [44], and the NHits framework has been applied to provide precise forecasts of COVID-19 incidence and mortality in Brazil [45]. In addition to these advancements, there has been a growing interest in modeling spatio-temporal dependencies in epidemic data. Traditional methods, such as the Generalized Spatiotemporal Autoregressive (GSTAR) [46] and Fast Gaussian Process (GpGp) [47] models, have been used to capture spatial and temporal patterns. Modern deep learning architectures, such as Spatiotemporal Graph Convolutional Networks (STGCN), have further enhanced the ability to model these complex relationships [48]. A key challenge remains to include the epidemiological knowledge and spatial dynamics of the disease inside the deep learning frameworks so that they can be used for informed decision-making in public health departments.
4 Modified Networked SIR Model
@@ 4.1 Model Formulation
where $y\sim x$ describes node $y$ is adjacent to node $x$ and $h$ is a function such that $h:V\to\mathbb{R}$, where $\mathbb{R}$ represents the set of real numbers [25, 27]. The operator $\Delta$ describes population mobility between regions. Building on this framework, combined with the model in Eq. (3), we develop the networked SIR epidemic model with a saturation incidence rate, incorporating graph Laplacian diffusion, as follows:
@@
$\dfrac{\partial R(x,t)}{\partial t}-\sigma\Delta R(x,t)$ $=\gamma I(x,t)-\mu R(x,t),\quad R(x,0)=R_{0}(x)\geq 0\mbox{ for }x\in V.$
The parameter $\sigma$, which ranges between 0 and 1, characterizes the rate of population mobility. This parameter is often referred to as the migration parameter or diffusion parameter. By adjusting $\sigma$, one can modulate population mobility as required. Notably, when $\sigma$ approaches 1, all regions tend to exhibit similar dynamics, behaving almost uniformly. In contrast, when $\sigma$ is near 0, the solution profiles appear decoupled, with regions behaving independently [49].
@@ 4.2 Equilibrium points
Disease-free equilibrium point is $(S_{0},0,0)$, where $S_{0}=\Lambda/\mu$ (notations have been discussed in Section 3.1) and endemic equilibrium point is $(S^{*},I^{*},R^{*})$ where
@@
Thus, the endemic equilibrium exists if $\mathcal{R}_{0}>1$, where the basic reproduction number ($\mathcal{R}_{0}$) is given by $\mathcal{R}_{0}=\dfrac{\beta\Lambda}{\mu(\gamma+\mu)}$. From an epidemiological perspective, $\mathcal{R}_{0}$ is the expected number of secondary cases generated by a typical infected individual throughout his infectious interval in a fully susceptible population. Mathematically, the progression of a disease is typically characterized by $\mathcal{R}_{0}$, such that if $\mathcal{R}_{0}<1$, the disease-free equilibrium is stable, and the disease will eventually die out. Conversely, if $\mathcal{R}_{0}>1$, the disease-free equilibrium becomes unstable, and the system transitions towards the endemic equilibrium, indicating the sustained presence of the disease in the population. Therefore, $\mathcal{R}_{0}$ determines the threshold quantity for investigating the asymptotic stability of the equilibrium states and the prediction value needed for disease eradication. To determine the equilibria states and analyze the stability of these states, we introduce the following results by [25]:
@@
where $a,\alpha$, and $c$ are positive constants. Also, note that Eq. (7) has a trivial equilibrium at $u_{0}=0$. If $a>c$, then the system in Eq. (7) has a unique positive non-trivial equilibrium, $u^{*}=\dfrac{a-c}{\alpha c}$.
@@
The proof of the lemma is given in Lemma 2.4 of [25].
@@ 4.3 Stability Analysis
To establish the global stability of an equilibrium point, we analyze the global dynamics of the disease-free equilibrium and the endemic equilibrium of the system Eq. (5) using Green’s formula and the Comparison principle. To simplify the analysis, we initially perform model dimension reduction. Since the variable $R$ does not appear in the first two equations of the system, the last equation is omitted to prove the global dynamics of the system in Eq. (5). Thus, we focus on the following subsystem:
@@
For the model to be epidemiologically meaningful, we show the positivity and boundedness of solutions for all time $t$ as follows:
@@
The disease-free equilibrium point denotes the point at which there are no infections in the population. On the contrary, the endemic equilibrium state denotes the state where the disease cannot be completely eradicated but persists in the population. To determine the stability of the disease-free equilibrium state and endemic equilibrium state, we have the following theorems:
@@
The global stability analysis ensures that, regardless of initial conditions, the disease will ultimately converge to one of the equilibrium states, such as the disease-free equilibrium or endemic equilibrium. This concept is central to understanding how epidemic models predict the long-term trajectory of disease spread under different intervention strategies. By examining these equilibria, we can assess the effectiveness of interventions in achieving disease control. Building on these epidemic insights, we integrate the epidemiological principles of the networked SIR model into deep learning techniques to generate reliable epidemic forecasts.
@@ 5 Epidemic-Guided Deep Learning Framework
This section outlines the generic architecture of the epidemic-guided deep learning (EGDL) methods, which seamlessly integrate epidemiological principles with spatial interactions modeled through the networked SIR framework (described in Section 4) with data-driven techniques to forecast the future trajectories of the disease incidence across different geographical locations. Specifically, we introduce two distinct approaches, namely EGDL-Parallel and EGDL-Series, which integrate the scientific knowledge of the disease dynamics and its spatial distribution into deep learning models to forecast TB incidence cases. The parameters of the networked SIR model, including transmission, recovery, saturation, population mobility, and birth/death rates, guide the deep learning architectures to ensure that the TB incidence forecasts align with real-world epidemiological trends and can effectively capture both the spatial and temporal dependencies. A pictorial representation (Fig A1) and pseudo-code (Algorithm 1) of the EGDL architectures are presented in Appendix B.
@@ 5.1 EGDL-Parallel Method
We consider the problem of forecasting epidemic incidence cases based on $T$ historical observations recorded across $n$ vulnerable locations. We denote the number of infected individuals observed at time $t$ in location $x$ as $Y(x,t);\;t=1,2,\ldots,T,\;x=1,2,\ldots,n$. Our goal is to iteratively generate the $q\;(\geq 1)$ step-ahead forecasts of $Y(x,t)$, to be denoted as $\widehat{Y}(x,{T+1}),\widehat{Y}(x,{T+2}),\ldots,\widehat{Y}(x,{T+q})$. A standard approach to this problem is to use a forecasting model $f_{M}\in\mathcal{F}$, where $\mathcal{F}$ represents the set of possible data-driven global forecasting models. In this study, we choose $\mathcal{F}$ to represent two types of data-driven deep learning architectures: stacked-based frameworks, such as NBeats [41] and NHits [42], and encoder-decoder-based architectures such as TCN [40] and Transformers [37]. Thus using the historical data $\{Y(1,t),Y(2,t),\ldots Y(n,t);t=1,2,\ldots,T\}$ for training $f_{M}$ the future dynamics can be predicted as follows:
where $x=1,2,\ldots,n,\;\tilde{t}=t_{w},t_{w}+1,\ldots,T,\;v=1,2,\ldots,q$, and $t_{w}$ denotes the number of historical inputs used for predicting the subsequent observations. Despite their effectiveness, the purely data-driven forecasting techniques $f_{M}$ solely rely on the lagged incidence time series without incorporating the scientific knowledge that governs the disease dynamics. As a result, this approach leads to a simple extrapolation of the temporal trends, which may fail to capture the underlying epidemiological underpinnings of the disease. To address this limitation, the EGDL-Parallel approach leverages the networked SIR model with saturated incidence rate and graph Laplacian diffusion to estimate the infection dynamics curve for the $n$ locations based on various disease drivers such as transmission, recovery, saturation, population mobility, and population death rates. The number of infection cases $I(x,t)$ predicted at timestamp $t$ for the $x^{th}$ location using the networked SIR infection curve captures the epidemic knowledge and models spatial dependencies among the neighboring locations using the graph Laplacian diffusion. However, these estimates provide an incomplete representation of the target variable $Y(x,t)$ due to their inability to track the lagged temporal dependencies, leading to discrepancies with ground truth observations.
The EGDL-Parallel approach overcomes this issue by integrating the networked SIR infection estimates $I(x,t)$ as auxiliary variables into the data-driven forecasting framework. This hybrid input allows the forecasting model to combine the strengths of both compartmental methods and data-driven techniques to forecast incidence cases. The resulting EGDL-Parallel framework, denoted as $f_{EGDLP}$, is defined as:
@@
where $f_{EGDLP}\in\mathcal{F}$ and notations are consistent with Eq. (12). The primary advantage of the $f_{EGDLP}$ model over its data-centric counterpart $f_{M}$ is its ability to integrate disease-specific knowledge, enhancing its generalizability and forecast accuracy. By leveraging the networked SIR model, the $f_{EGDLP}$ framework is guided by epidemiological principles, enabling the generation of realistic future estimates of disease incidence cases. Moreover, it effectively learns the complex spatiotemporal features from the compartmental models and data-driven methods, making it a robust solution for epidemic forecasting.
@@ 5.2 EGDL-Series Method
The EGDL-Series framework employs an integrated deterministic-stochastic approach to model and forecast disease incidence cases by combining epidemiological principles with data-driven techniques. In the EGDL-Series method, a deterministic model (networked SIR in this case) is combined with stochastic models such as NBeats [41], NHits [42], Transformers [37], and TCN [40] using a residual remodeling technique [51, 52]. Mathematically, the disease incidence cases recorded at time $t$ for a location $x$ can be expressed as: $Y(x,t)=\mathcal{D}(x,t)+\mathcal{\tilde{S}}(x,t),$ where $\mathcal{D}(x,t)$ represent the deterministic component and $\mathcal{\tilde{S}}(x,t)$ accounts for the stochastic variations in $Y(x,t)$. To forecast future disease trajectories, the EGDL-Series framework first utilizes the networked SIR model with saturated incidence rate and graph Laplacian diffusion to estimate the infection dynamics $I(x,t)$. This curve, determined based on key epidemiological factors, such as disease transmission, recovery, saturation, population mobility, and birth/death rates, approximates the deterministic components $\mathcal{D}(x,t)$ of $Y(x,t)$. The predicted infection trajectory, $\widehat{\mathcal{D}}(x,t)\equiv I(x,t)$, captures structural variations in the disease incidence by modeling epidemic principles and spatial interactions among neighboring locations. The deviations between observed incidences and the predictions of the networked SIR model, defined as residuals $e(x,t)=Y(x,t)-I(x,t)$, capture the stochastic variations that the deterministic part fails to account for. These residuals store nonlinearities, temporal dependencies, and unexplained stochastic dynamics that the networked SIR framework cannot model. The residual series is modeled using data-driven forecasting techniques, denoted as $f_{EGDLS}\in\mathcal{F}$ to address this. The $q$-step-ahead forecasts of the residual series generated by the deep learning-based temporal forecasters are combined with the predictions of the networked SIR model to forecast the disease incidence as follows:
$\widehat{Y}(x,\tilde{t}+v)=I(x,\tilde{t}+v)+\hat{e}(x,\tilde{t}+v),$
with $\hat{e}(x,\tilde{t}+v)=f_{EGDLS}\left(\left\{e\left(x,\tilde{t}-t_{0}\right)\right\}_{t_{0}=0}^{t_{w}-1}\right),$ notations are consistent with Eq. (12). The proposed formulation ensures that the model aligns with the epidemic principles coupled with spatial interactions while simultaneously addressing the unexplained variations of the compartmental model. The EGDL-Series approach offers enhanced flexibility, greater insight into disease dynamics, and improved forecasting accuracy by integrating compartmental and data-driven components.

### v2

@@ 3.3 Deep Learning Models for Time Series Forecasting
Time series forecasting has become a crucial area of research in public health, particularly for modeling the spread and impact of infectious diseases. The increasing availability of epidemic-related data has supported the development of data-driven forecasting methods aimed at early detection of abnormal disease patterns and mitigation of public health risks [39, 40, 41]. These methods are essential for timely interventions and effective surveillance systems. Recently, deep learning approaches have emerged as powerful tools due to their ability to capture complex, non-linear dependencies in high-dimensional time series data, often outperforming traditional statistical models [17]. Architectural innovations in deep learning-based forecasting have taken various forms. Encoder-decoder models with attention mechanisms, most notably Transformers, have proven effective in learning long-range temporal dependencies [42, 43]. These models process input sequences using positional encodings and stacked self-attention layers, followed by decoder components that leverage look-ahead masking to ensure predictions depend only on past observations. Similarly, Temporal Convolutional Networks (TCNs) offer an alternative by using dilated causal convolutions within residual blocks to extract temporal features and model uncertainty in sequential data [44]. On the other hand, attention-free and convolution-free models with deep stacks of fully connected layers have also shown strong performance in capturing input-output relationships while implicitly prioritizing resource-efficient computations. Among these models, the Neural Basis Expansion for Time Series (NBeats) framework uses a block-based architecture to forecast time series data through residual learning [45]. It consists of stacked blocks, where the initial block models the input time series and generates forward and backward outputs. Each subsequent block refines the residuals of the previous one. Within each block, dense layers with ReLU activations produce expansion coefficients, which are then passed through backward and forward basis layers to generate the backcast and forecast. Building upon this architecture, the Neural Hierarchical Interpolation for Time Series (NHits) framework introduces a multi-rate signal sampling approach to improve long-range forecasting [46]. Each NHits block comprises a Multilayer Perceptron (MLP) followed by MaxPool operations to focus on analyzing low-frequency components crucial for stable long-range forecasting.
@@
The architectural designs of these scalable deep learning forecasters have demonstrated superior performance in capturing long-term dependencies of epidemic datasets. For instance, Transformers have been shown to accurately forecast influenza cases [43, 47], while attention mechanisms have been utilized for predicting dengue incidence in Vietnam [48]. Similarly, a modified NBeats framework was developed to forecast COVID-19 hospitalizations in Canada [49], and the NHits framework has been applied to provide precise forecasts of COVID-19 incidence and mortality in Brazil [50]. In addition to these advancements, there has been a growing interest in modeling spatio-temporal dependencies in epidemic data. Traditional methods, such as the Generalized Spatiotemporal Autoregressive (GSTAR) [51] and Fast Gaussian Process (GpGp) [52] models, have been used to capture spatial and temporal patterns. Modern deep learning architectures, such as Spatiotemporal Graph Convolutional Networks (STGCN), have further enhanced the ability to model these complex relationships [53]. A key challenge remains to include the epidemiological knowledge and spatial dynamics of the disease inside the deep learning frameworks so that they can be used for informed decision-making in public health departments.
4 Modified Networked SIR (MN-SIR) Model
@@ 4.1 MN-SIR Model Formulation
where $y\sim x$ describes node $y$ is adjacent to node $x$ and $h$ is a function such that $h:V\to\mathbb{R}$, where $\mathbb{R}$ represents the set of real numbers [27, 29]. The operator $\Delta$ describes population mobility between regions. Building on this framework, combined with the model in Eq. (3), we develop the MN-SIR epidemic model with a saturation incidence rate, incorporating graph Laplacian diffusion, as follows:
@@
$\dfrac{\partial R(x,t)}{\partial t}-\sigma\Delta R(x,t)$ $=\gamma I(x,t)-\mu R(x,t),\quad R(x,0)=R_{0}(x)\geq 0\mbox{ for }x\in V,$
where $S_{0}(x),I_{0}(x),$ and $R_{0}(x)$ denote the initial values of the susceptible, infected, and recovered populations, respectively, at node $x\in V$. Model (6) captures the spatio-temporal dynamics of the susceptible ($S(x,t)$), infected ($I(x,t)$), and recovered ($R(x,t)$) populations across space ($x\in V$) and time ($t$). It combines local disease transmission processes with spatial diffusion, representing the movement or interaction of individuals across space. The partial derivatives $\frac{\partial S(x,t)}{\partial t}$, $\frac{\partial I(x,t)}{\partial t}$, and $\frac{\partial R(x,t)}{\partial t}$ represent the temporal rates of change for each compartment at location $x$. The diffusion terms $\sigma\Delta S(x,t)$, $\sigma\Delta I(x,t)$, and $\sigma\Delta R(x,t)$ describe the spatial movement of individuals, where $\sigma$ is the diffusion coefficient or migration parameter, ranging between 0 and 1. Adjusting $\sigma$ governs the rate of population mobility across locations, such that $\sigma\to 1$ indicates all regions exhibit nearly uniform dynamics, while $\sigma\to 0$ leads to decoupled, independent regional behavior [54]. The nonlinear transmission term $\frac{\beta S(x,t)I(x,t)}{1+\alpha I(x,t)}$ introduces a saturation effect, reflecting the slowing growth of transmission when the infected population becomes large. This captures the realistic scenario where transmission does not increase indefinitely but saturates due to factors such as limited contact opportunities, improved healthcare access, increased public awareness, among many others.
@@ 4.2 Equilibrium points
Disease-free equilibrium point is $(S_{0},0,0)$, where $S_{0}=\Lambda/\mu$ (notations have been discussed in Section 3.1) and endemic equilibrium point is $(S^{*},I^{*},R^{*})$, where
@@
Thus, the endemic equilibrium exists if $\mathcal{R}_{0}>1$, where the basic reproduction number ($\mathcal{R}_{0}$) is given by $\mathcal{R}_{0}=\dfrac{\beta\Lambda}{\mu(\gamma+\mu)}$. From an epidemiological perspective, $\mathcal{R}_{0}$ is the expected number of secondary cases generated by a typical infected individual throughout his infectious interval in a fully susceptible population. Mathematically, the progression of a disease is typically characterized by $\mathcal{R}_{0}$, such that if $\mathcal{R}_{0}<1$, the disease-free equilibrium is stable, and the disease will eventually die out. Conversely, if $\mathcal{R}_{0}>1$, the disease-free equilibrium becomes unstable, and the system transitions towards the endemic equilibrium, indicating the sustained presence of the disease in the population. Therefore, $\mathcal{R}_{0}$ determines the threshold quantity for investigating the asymptotic stability of the equilibrium states and the prediction value needed for disease eradication. To determine the equilibria states and analyze the stability of these states, we introduce the following Lemma by [27], which will be utilized in the proof of stability analysis:
@@
where $a,\alpha$, and $c$ are positive constants, and $\dot{u}$ represents the derivative of $u$ with respect to time. Also, note that Eq. (8) has a trivial equilibrium at $u_{0}=0$. If $a>c$, then the system in Eq. (8) has a unique positive non-trivial equilibrium, $u^{*}=\dfrac{a-c}{\alpha c}$.
@@
The proof of the lemma is given in Lemma 2.4 of [27]. We present the above results to facilitate the proof of Lemma 4.4.
@@ 4.3 Stability Analysis
To establish the global stability of an equilibrium point, we analyze the global dynamics of the disease-free equilibrium and the endemic equilibrium of the system in Eq. (6) using Green’s formula and the Comparison principle. To simplify the analysis, we initially perform model dimension reduction. Since the variable $R$ does not appear in the first two equations of the system, the last equation is omitted to prove the global dynamics of the system in Eq. (6). Thus, we focus on the following subsystem:
@@
For the model to be epidemiologically meaningful, we show the positivity and boundedness of solutions for all time $t$, since the population sizes cannot be negative or unbounded.
@@
The disease-free equilibrium point denotes the point at which there are no infections in the population. On contrary, the endemic equilibrium state denotes the state where the disease cannot be completely eradicated but persists in the population. To determine the stability of the disease-free equilibrium state and endemic equilibrium state, we have the following theorems:
@@
The global stability analysis ensures that, regardless of initial conditions, the disease will ultimately converge to one of the equilibrium states, such as the disease-free equilibrium or endemic equilibrium. This concept is central to understanding how epidemic models predict the long-term trajectory of disease spread under different intervention strategies. By examining these equilibria, we can assess the effectiveness of interventions in achieving disease control. Building on these epidemic insights, we integrate the epidemiological principles of the MN-SIR model into deep learning techniques to generate reliable epidemic forecasts.
@@ 5 Epidemic-Guided Deep Learning Framework
This section outlines the generic architecture of the epidemic-guided deep learning (EGDL) methods, which seamlessly integrate epidemiological principles with spatial interactions modeled through the MN-SIR framework (described in Section 4) with data-driven techniques to forecast the future trajectories of the disease incidence across different geographical locations. Specifically, we introduce two distinct approaches, namely EGDL-Parallel and EGDL-Series, which integrate the scientific knowledge of the disease dynamics and its spatial distribution into deep learning models to forecast TB incidence cases. The parameters of the MN-SIR model, including transmission, recovery, saturation, population mobility, and birth/death rates, guide the deep learning architectures to ensure that the TB incidence forecasts align with real-world epidemiological trends and can effectively capture both the spatial and temporal dependencies. Fig. 3 illustrates the schematic architecture of the EGDL frameworks. A detailed pseudo-code for the proposed approaches is presented in Algorithm 1 of Appendix B.
Figure 3: Epidemic-Guided Deep Learning (EGDL). The upper panel of the plot represents a schematic architecture of the modified networked SIR (MN-SIR) model, and the lower panel showcases the workflow of the EGDL-Parallel framework (left) and the EGDL-Series approach (right). The demonstration is given using Japan’s TB dataset.
@@ 5.1 EGDL-Parallel Method
We consider the problem of forecasting epidemic incidence cases based on $T$ historical observations recorded across $n$ vulnerable locations. Following the notations introduced in Section 3.2, the objective is to iteratively generate the $q$-step-ahead forecasts of $\{Y(x,t)\}_{t\leq T}$ as $\{\widehat{Y}(x,T+v)\}_{v=1}^{q}$ for each location $x\in V$. A standard approach is to use a global forecasting model $f_{M}\in\mathcal{F}$, as described in Eq. (4), which uses past incidence data to generate future forecasts. In this study, we choose $\mathcal{F}$ to represent two broad categories of data-driven deep learning architectures: stacked-based frameworks such as NBeats and NHits, and encoder-decoder-based models such as TCN and Transformers. While these models are effective at capturing temporal dependencies and complex nonlinear patterns, they solely rely on the lagged observations of the incidence time series without incorporating the underlying epidemiological mechanisms governing disease spread. As a result, this approach leads to a simple extrapolation of the temporal trends, which may fail to capture the underlying epidemiological underpinnings of the disease. To address this limitation, the EGDL-Parallel approach leverages the MN-SIR model to estimate the infection dynamics curve for the $n$ locations based on various disease drivers such as transmission, recovery, saturation, population mobility, and population death rates. The number of infection cases $I(x,t)$ predicted at timestamp $t$ for location $x$ using the MN-SIR model’s infection curve captures the epidemic knowledge and models spatial dependencies among the neighboring locations using the graph Laplacian diffusion. However, these estimates provide an incomplete representation of the target variable $Y(x,t)$ due to their inability to track the lagged temporal dependencies, leading to discrepancies with ground truth observations.
The EGDL-Parallel approach overcomes this issue by integrating the MN-SIR infection estimates $I(x,t)$ as auxiliary variables (exogenous information) into the data-driven forecasting framework. This hybrid input allows the forecasting model to combine the strengths of both compartmental methods and data-driven techniques to forecast incidence cases. The resulting EGDL-Parallel framework, denoted as $f_{EGDLP}$, is defined as:
@@
where $f_{EGDLP}\in\mathcal{F}$ and notations are consistent with Eq. (4). The primary advantage of the $f_{EGDLP}$ model over its data-centric counterpart $f_{M}$ is its ability to integrate disease-specific knowledge, enhancing its generalizability and forecast accuracy. By leveraging the MN-SIR model, the $f_{EGDLP}$ framework is guided by epidemiological principles, enabling the generation of realistic future estimates of disease incidence cases. Moreover, it effectively learns the complex spatiotemporal features from the compartmental models and data-driven methods, making it a robust solution for epidemic forecasting.
@@ 5.2 EGDL-Series Method
The EGDL-Series framework employs an integrated deterministic-stochastic approach to model and forecast disease incidence cases by combining epidemiological principles with data-driven techniques. In the EGDL-Series method, a deterministic model (MN-SIR in this case) is combined with stochastic models from set $\mathcal{F}$ using a residual remodeling approach [56, 57]. Mathematically, the disease incidence cases recorded at time $t$ for a location $x$ can be expressed as: $Y(x,t)=\mathcal{D}(x,t)+\mathcal{\tilde{S}}(x,t),$ where $\mathcal{D}(x,t)$ represent the deterministic component and $\mathcal{\tilde{S}}(x,t)$ accounts for the stochastic variations in $Y(x,t)$. To forecast future disease trajectories, the EGDL-Series framework first utilizes the MN-SIR model to estimate the infection dynamics $I(x,t)$. This curve, determined based on key epidemiological factors, such as disease transmission, recovery, saturation, population mobility, and birth/death rates, approximates the deterministic components $\mathcal{D}(x,t)$ of $Y(x,t)$. The predicted infection trajectory, $\widehat{\mathcal{D}}(x,t)\equiv I(x,t)$, captures structural variations in the disease incidence by modeling epidemic principles and spatial interactions among neighboring locations. The deviations between observed incidences and the predictions of the MN-SIR model, defined as residuals $e(x,t)=Y(x,t)-I(x,t)$, capture the stochastic variations that the deterministic part fails to account for. These residuals store nonlinearities, temporal dependencies, and unexplained stochastic dynamics that the MN-SIR framework cannot model. The residual series is modeled using data-driven forecasting techniques, $f_{M}\in\mathcal{F}$. The $q$-step-ahead forecasts of the residual series generated by the deep learning-based temporal forecasters are combined with the predictions of the MN-SIR model to forecast the disease incidence. The predictions of the EGDL-Series methods can be mathematically represented as:
$\widehat{Y}(x,\tilde{t}+v)=\hat{f}_{EGDLS}(x,\tilde{t}+v)=I(x,\tilde{t}+v)+\hat{e}(x,\tilde{t}+v),$
with $\hat{e}(x,\tilde{t}+v)=f_{M}\left(\left\{e\left(x,\tilde{t}-t_{0}\right)\right\}_{t_{0}=0}^{t_{w}-1}\right),f_{M}\in\mathcal{F},$ following the same notations as in Eq. (4). The proposed formulation ensures that the model aligns with the epidemic principles coupled with spatial interactions while simultaneously addressing the unexplained variations of the compartmental model. The EGDL-Series approach offers enhanced flexibility, greater insight into disease dynamics, and improved forecasting accuracy by integrating compartmental and data-driven components.